OpenStudy (anonymous):

The limit represents the derivative of some function f at some number a. State such an f and a. lim h--->0 ((1+h)^7-1)/h

OpenStudy (anonymous):

From the definition of the derivative: \[f'(a)=\lim_{h\to0}\frac{f(a+h)-f(a)}{h}\] You want to find a function \(f\) such that \[\begin{cases}f(a+h)=\dfrac{1}{(1+h)^7}\\\\ f(a)=1\end{cases}\] Hint: \[f(a)=1=\frac{1}{1^7}\]

OpenStudy (anonymous):

f(x) = x^6, a = 2 f(x) = x^7, a = 1 f(x) = x^8, a = 0 f(x) = x^7 − x, a = 1 f(x) = x^7 + x, a = 0 these are the answer choices and i know the third one is wrong but im guessing it could be the second one? @SithsAndGiggles

OpenStudy (anonymous):

the fourth one is supposed to say f(x)=x^7-x, a=1

OpenStudy (anonymous):

Oh I must have misread, you have \[\begin{cases}f(a+h)=(1+h)^7\\ f(a)=1\end{cases}\] And you have \[f(a)=1=1^7\] Suppose \(f(x)=x^7\). Then indeed, \(f(1+h)=(1+h)^7\), and if \(a=1\) then \(f(a)=1^7=1\).

OpenStudy (anonymous):

thank you, i understand it now! @SithsAndGiggles